OpenStudy (anonymous):

Find the least common multiple of the set of numbers 3, 30, 12, 8

OpenStudy (anonymous):

Do you know the formula for the least common multiple?

OpenStudy (anonymous):

I've done this problem at least five times.. (With different numbers because everytime I get it wrong it resets) but @tom982 Obviously not..

OpenStudy (anonymous):

@Markiplier, that's completely wrong. @kaela_stanley, the least common multiple of a set of integers is the product of these integers divided by the highest common factor (also known as the greatest common divisor). Can you tell me what the highest common factor of these numbers is?

OpenStudy (anonymous):

@tom982 Would that be 2?

OpenStudy (anonymous):

Spot on, great work! Now for the LCM, can you tell me what that is?

OpenStudy (anonymous):

3,1,5,2,2,2 ? Which would = 120? @tom982

OpenStudy (anonymous):

That's the answer, but you got there the wrong way (I think it's fluke that you were correct) which is partly my fault - I forgot to explain one part, sorry! We know LCM(a,b)=(a*b)/HCF(a,b) but we have more than two numbers here, so we just pick two to start off with from our set: 3, 30, 12, 8. I'll pick the first two: GCD(3,30)=3 HCF(3,30)=3*30/3=30 Replace 3 and 30 with the HCF: 30. Our set is now 30, 12, 8. GCD(30, 12)=6 HCF(30,12)=30*12/6=60 Replace 30 and 12 with the HCF: 60. Our set is now 60 and 8. GCD(60,8)=4 HCF(60,8)=60*8/4=120 And that's our answer :)